OpenStudy (aravindg):

Show that the function defined by f:QQ:f(x)=3x + 5 for all x Є Q is one-one and onto. Find a formula for f -1.

OpenStudy (anonymous):

last question first. this function says 1) multiply by 3 2) add 5 inverse says 1) subtract 5 2) divide by 3 \[f^{-1}(x)=\frac{x-5}{3}\]

OpenStudy (aravindg):

k

OpenStudy (anonymous):

to show it is one to one you have a choice. you can say " it is a line and all lines are one to one (except horizontal ones) because they all pass the horizontal line test"

OpenStudy (anonymous):

or you can say Let \[f(a)=f(b)\] and show what \[a=b\] as follows \[f(a)=f(b)=\] \[3a+5=ab+5\iff 3a=ab\iff a=b\]

OpenStudy (aravindg):

wow

OpenStudy (aravindg):

u r a legend indeed

OpenStudy (anonymous):

you also have a simple way to find the inverse if you do did not like my words. write \[y=3x+5\] then switch x and y to write \[x=3y+5\] solve for y get \[x-5=3y\] \[\frac{x-5}{3}=y\] so \[f^{-1}(x)=\frac{x-5}{3}\] as before

OpenStudy (aravindg):

how do we say its onto???/

OpenStudy (aravindg):

u there???

OpenStudy (anonymous):

onto?

OpenStudy (aravindg):

ya

OpenStudy (anonymous):

show that for any number "b" there is an "a" with \[f(a)=b\] but you have already done this because you know if that \[a=\frac{b-5}{3}\]

OpenStudy (aravindg):

wow thx a lott